i'm suppose to solve it or simplify it if it is an equation

17) 8 + SR(-20) divided by -4

128) 6a^2 - 4 = -8

129) 3 + SR(-9) =

Answer by stanbon(75887)

17) 8 + SR(-20) divided by -4
[8 + 2isqrt(5)]/-4
= [-2 - (1/2)sqrt(5)]
======================
128) 6a^2 - 4 = -8
6a^2 +4 = 0
a^2 = -2/3
a = +-i[sqrt(6)/3]
======================

129) 3 + SR(-9) =
= 3 + 3i
======================
Cheers,
Stan H.